You would like to buy a house, but you can only afford monthly payments of R500. Thirty-year loans are available at 6% interest compounded monthly. If you can make a R5 000 down payment, what is the price of the most expensive house that you can afford to purchase?


1
Expert's answer
2022-04-04T09:15:24-0400

"P=M*{(1+\frac{r}{12})^n-1 \over (\frac{r}{12})(1+{r \over 12})^n}+DownPayment"

"P=500*{(1+\frac{0.06}{12})^{360}-1 \over (\frac{0.06}{12})(1+{0.06 \over 12})^{360}}+5000=88,395.81"